how many lattice points are their in one unit cell of each of the following lattice?
1)face centered cubic
2)face centered tetragonal
3)body centered

Grade:12

5 Answers

BOLLU SRINIVAS
8 years ago
(i) Number of corner atoms per unit cell= 8 corners atoms X1/8 atom per unit cell= 8 X 1/8 = 1 atomNumber of face centred atoms per unit cell= 6 face centred atoms X 1/2 atom per unit cell= 6 X1/2 = 3 atoms
BOLLU SRINIVAS
8 years ago
 Total number of atoms or lattice points = 1 + 3 = 4.
(ii) No. of lattice points in Face-centred tetragonal = 4.
(iii) In bcc unit cell, number of corner atoms per unit cell
= 8 corners X 1/8 per corner atom.
= 8 X 1/8 = 1 atom
Number of atoms at body centre = 1 X 1 = 1 atom
Total number of atoms (lattice points) = 1 + 1 = 2.
 
Priyanshu
5 years ago
Lattice point are there in 1 unit cell in each of the following letters are as follows:-(1). Face centered cubic have 4(2). Face centred tetragonal have 4(3). Body centred have 2
Gitanjali Rout
5 years ago
i) Number of corner atoms per unit cell= 8 corners atoms X1/8 atom per unit cell= 8 X 1/8 = 1 atomNumber of face centred atoms per unit cell= 6 face centred atoms X 1/2 atom per unit cell= 6 X1/2 = 3 atoms
Gitanjali Rout
5 years ago
otal number of atoms or lattice points = 1 + 3 = 4.
(ii) No. of lattice points in Face-centred tetragonal = 4.
(iii) In bcc unit cell, number of corner atoms per unit cell
= 8 corners X 1/8 per corner atom.
= 8 X 1/8 = 1 atom
Number of atoms at body centre = 1 X 1 = 1 atom